OpenStudy (anonymous):

Evaluate the integral of 1/(x^2-9)^(3/2)dx Please help!

OpenStudy (anonymous):

Umm, use a trig sub.... \(x= 3\sec(\theta)\).

OpenStudy (anonymous):

dx=what?

OpenStudy (anonymous):

\[ dx =\frac{dx}{d\theta}d\theta= 3\sec(\theta)\tan(\theta) d\theta \]You don't remember how to do your derivatives?

OpenStudy (anonymous):

Its been a long time since I've covered this stuff...Thats why I am here. :P

OpenStudy (anonymous):

So how do you plan on being able to replace the original problem with that x and dx?

OpenStudy (anonymous):

if you don't mind me mentioning, but I believe if you take a look at Paul's website at http://tutorial.math.lamar.edu/Classes/CalcII/TrigSubstitutions.aspx He covers the basic trig substitution methods there. I believe this will give you a better insight of how and especially why to approach such problems in the suggested way. Since you suggested yourself that you're reviewing this stuff, I think this will make things much clearer.

OpenStudy (anonymous):

\[ \int \frac{1}{(x^2-9)^{3/2}}dx = \int \frac{3\sec\theta\tan\theta}{((3\sec\theta)^2-9)^{3/2}}d\theta \]

OpenStudy (anonymous):

Okay so what next?

OpenStudy (anonymous):

Now I set u= sec(theta)

OpenStudy (anonymous):

\[ \sec^2(\theta) - 1 = \tan^2(\theta) \]

OpenStudy (anonymous):

So the bottom is like ((3tan^2(theta)-9)^3/2?

OpenStudy (anonymous):

\[ (3\sec\theta)^2-9 = 9\sec^2\theta-9 = 9(\sec^2\theta-1)=9\tan^2\theta = (3\tan\theta)^2 \]

OpenStudy (anonymous):

The bottom ends up becoming: \[ 3^3|\tan\theta|^3 \]

OpenStudy (anonymous):

Okay...Thanks..So as we continue what happens to the top?